A ball is thrown vertically upwards it was observed twice at a height of 'h' and a time interval of 't'.calculate initial velocity of body in terms of 'g' 't' and 'h'

Grade:

2 Answers

Abhishree Chowdhary
11 years ago

ANS=1/2root(g^2*t^2-8gh) .......?

if correct then please aproove ill post the solution

dhaval singh
11 years ago

√take the up direction as +ve. here s=h.

 substituting the value values in s=ut+1/2at2

h=ut-gt2

gt2-ut+h=0

t=u+√u2-4gh/2g    and   u-√u2-4gh/2g 

now the time is given t...

u+√u2-4gh/2g -   u-√u2-4gh/2g =t

solving this u2=g2t2+4gh

therefore u=√g2t2+4gh.